# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ I.4 Ion trap arrays and optical dipole traps
In this review we focus on ion crystals that form at low temperatures in a configuration resulting from the interplay of a common ion trap potential and the Coulomb repulsion. In this section we provide a short outlook on other kinds of trapping potentials that have been discussed in the literature and are based on different mechanisms than the ones discussed in the rest of this review. One such alternative approach consists of using an array of trapping potentials, where each cell stores a single ion [285]. This approach enables the generation of arbitrary lattice geometries. Ion trap arrays are being pursued both with rf ion traps [212, 199] as well as Penning traps [156, 157]. To date the number of ions stored in an ion trap array is small $(N\lesssim 3$ or $4)$.
A further approach employs optical dipole traps for ions with optical dipole transitions. A major motivation is to avoid boundary effects that are significant in Penning and Paul traps, thereby ideally realizing strings of equidistant ions. The micromotion due to the coupling of the charge monopole to the fast laser oscillations at the optical frequency is negligible [64], and the potential the ions experience is essentially the same as the potential experienced by neutral atom scattering in the visible. Optical trapping ion Coulomb crystals is challenging because of the strength of the Coulomb repulsion of closely spaced ions, requiring large laser powers. Optical trapping of ion crystals include trapping linear strings up to 6 ions by a single beam optical dipole trap [282], and confinement of a few ions in a one-dimensional optical lattice [141].
Somewhat less demanding in laser power is the use of optical dipole forces to modify the normal mode spectrum of trapped ion crystals. We refer the reader to Sec. III.5, where these experiments are extensively discussed.
@@ II Ion Coulomb crystals: from one to three dimensions
Trapped ions can crystallize in different structures, determined by the geometry of the confining potential and by their number. Linear, planar, up to cubic-centered structures have been realized in the laboratory. In this section we review the equilibrium properties of the crystals from one to three dimensional spatial arrangements of the ions. We start with the linear chain, on which an extensive number of studies have been performed, review then planar structures, and finally report on the crystalline properties of three dimensional structures. Differing from previous reviews, we set the focus on the dynamical and thermodynamical properties, for the purpose of establishing a direct link with condensed-matter systems.
@@ II.1 Linear geometries: the ion chain
Chains of ions are a laboratory realization of a one-dimensional Debye crystal [7]. They typically form at Doppler cooling temperatures. The interparticle distances are of the order of ten micrometers, and the individual ions can be imaged on a CCD camera via fluorescence light emitted during laser excitation. First experimental demonstrations were in quadrupole traps [22, 260, 328]. In the meantime linear chains have also been realized in Penning traps [203] and ring geometries [191]. Figure 8 is the image of an experimentally realized ion chain composed by more than 100 of ions [242]. The ion chain has been the workhorse of the ion-trap quantum computer for the versatility in controlling the individual components, such as the interactions between the electronic and the vibrational degrees of freedom at the quantum level [59]. Due to the Coulomb interactions, it constitutes a peculiar one-dimensional structure. Below we review salient dynamic and thermodynamic properties, emphasizing analogies with and fundamental differences from textbook examples of condensed matter systems.
Figure 8: (color online) Linear chain of 121 171Yb+ ions in a cryogenic linear Paul trap with $\omega_{y}/2\pi=1.5$ MHz and $\omega_{x}/2\pi=35$ kHz. The red arrow marks a lone 171Yb+ ion which is in a state that does not fluoresce. This dark ion retained its position after more than 4 hours indicating that the chain never melted despite collisions with the background gas. From 242.
@@ II.1.1 The ground state configuration
where $m$ is the mass, $Q$ the charge, and $\nu$ and $\nu_{t}$ are respectively the trap axial frequency and the transverse trap frequencies in the pseudopotential approximation, such that $\nu<\nu_{t}$. The linear chain is stable for $\nu_{t}/\nu$ exceeding a critical value, at which it undergoes a second order phase transition to a so-called zigzag array.
@@
with $|x|\leq L$ and vanishes otherwise. The parameter $L$ is the chain half-length at equilibrium and is determined by means of a variational ansatz, such that
$L(N)=\left[N\log N(3Q^{2}/(m\nu^{2})\right]^{1/3}\,$, (29)
to leading order in an expansion in $\log N$ [95]. Figure 9 shows that expression (28) gives a good estimate of the charge distribution sufficiently far away from the chain edges. At the chain center the density varies very slowly about the value $n(0)\sim N/L$, where the interparticle distance is minimal, $a(0)=1/n(0)$.
Figure 9: Density of charges per unit length, $n(x)$, as a function of position $x$ from the center of the chain for different numbers $N$ of ions in a linear Paul trap. The dots are the equilibrium positions determined by molecular dynamics simulation for a finite number of ions $N$. The lines are obtained using the method of trial variational functions with two parameters. The dashed lines are the one-parameter function of Eq. (28). Both $x$ and $n(x)$ are scaled to $(Q^{2}/m\nu^{2})^{1/3}$. From 95.
@@
Strictly speaking, thermal and quantum fluctuations destroy diagonal long-range order in one dimension, thus the linear chain is unstable in the thermodynamic limit. Yet, this is an academic argument. Stability against quantum fluctuations has been studied in [287] by means of a treament based on bosonization, showing that at zero temperature the density-density correlations of a one-dimensional Wigner crystal decay with the distance $x$ as $\exp(-c\sqrt{\log x})$ with $c$ a constant [287]. This is an extremely slow decay: any large (but finite) chain can be considered crystallized. Stability against thermal fluctuations requires $Q^{2}\log N/(4\pi\epsilon_{0}a_{0})\gg\kappa_{B}T$, namely, the potential energy scale is larger than the thermal fluctuations where $T$ is the temperature and the classical equipartition theorem has been used [224]. This shows that, for sufficiently large chains, the long-range Coulomb correlations dominate over thermal disorder.
II.1.2 Normal modes of the ion chain in a linear Paul trap
Figure 10: The frequency spectrum of a laser cooled chain composed of 24 9Be+ ions. Dashed lines are the experimental results of normalized excitation level versus frequency, where red, black, and blue denote the scan over red sideband (anti-Stokes), carrier, and blue sideband (Stokes) transitions, respectively. The brightness of the dashed line shows the variable probe times throughout the spectrum to ensure that all modes are excited with high visibilities. Dashed vertical blue and red lines show all 24 calculated first-order red and blue sideband frequencies. Dashed vertical green lines show the mode mixings between c.m. and a few higher frequency modes, coinciding with the smaller peaks in the experimental data. Details on the figures can be found in Ref. 335.
At laser cooling temperatures, the ions perform vibrations about the equilibrium positions with displacements of the order of tens of nanometers, and thus about three orders of magnitude smaller than the equilibrium distance. In this regime the restoring forces are harmonic and the ion chain is a one-dimensional harmonic crystal. Figure 10 displays the Stokes sidebands, measured in the resonance fluorescence of a chain of 24 ions. The measured spectra are reproduced by theoretical models that approximate the potential of Eq. (27) by only keeping the quadratic terms in the displacements about the equilibrium position. Below we focus on ion chains in a linear Paul trap.We refer to Refs. [195, 143, 48] for studies of the normal modes of ion chains in anharmonic traps.
@@ II.1.2 Normal modes of the ion chain in a linear Paul trap
with $\delta_{ij}$ Kronecker’s delta. Here, $\xi_{i}=x_{i}^{(0)}/L$ is the dimensionless equilibrium position rescaled by the half-length $L$ of the chain at equilibrium. The set of differential equations (32) and (33) is solved in Fourier space by diagonalizing the matrix $I_{ij}$, thus solving the linear set of equations
@@
$\omega^{\perp\,2}_{n}=\nu_{t}^{2}-K_{0}\lambda_{n}/2\,$, (37)
@@
In the rest of this section we make general considerations on the structure of the eigenvalues for $\nu_{t}>\nu_{t}^{(c)}$, which are based on the treatment of Ref. [225, 224]. For convenience, we adopt the convention $\lambda_{n}<\lambda_{m}$ for $n<m$. With this convention the label $n$ directly gives the number of nodes $\ell=n-1$ of the corresponding eigenmode $w^{(n)}$. Since the position of the nodes is generally constrained to the position of the ions within the chain, the value of $\lambda_{n}$ generally depends on the number of ions $N$. There are two exceptions: (i) The eigenvalue $\lambda_{1}=0$, corresponding to the bulk eigenmodes where the chain oscillates as a rigid body at the frequencies of the trap, and (ii) the eigenvalue $\lambda_{2}=2\nu^{2}/K_{0}$. These two eigenvalues are related to two symmetries of the Hamiltonian: point (i) is due to the fact that the dynamics is separable into center-of-mass and relative coordinates even in the presence of the harmonic trap, point (ii) is a consequence of the symmetry by reflection about the trap center. The breathing oscillations $\omega_{2}^{\|}=\sqrt{3}\nu$ and $\omega_{2}^{\perp}=\sqrt{\nu_{t}^{2}-\nu^{2}}$, in fact, are also characteristic of a one-component plasma in an ellipsoidal trap [99].
Figure 11: (a) Transverse and (b) axial spectrum of eigenfrequencies (in units of $\nu$) for a chain of $N=1000$ ions and with $\nu_{t}=400\nu$. The solid line is obtained from the numerical solution of Eq. (35), the grey line from the spectrum of Eq. (38) (these curves have been truncated, as they do not correctly reproduce the shortwavelength eigenmodes). The dotted line gives the spectra evaluated extending a method first developed for determining the density of states of a disordered chain. Adapted from Ref. 224.
@@
The eigenmodes are Jacobi polynomials $w^{(n)}=P_{\ell}^{(1,1)}$ with $n=\ell+1$ and $\ell=0,1,\ldots$. The eigenvalues read ${\lambda}_{n}=\ell(\ell+3)$ and exactly reproduce the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ for $\ell=0,1$, as also visible in Fig. 11. Interestingly, the model of a plasma in a cigar-shaped trap potential gives the same expression for the axial spectrum $\omega_{n}^{\|}=\nu\sqrt{n(n+1)/2}$ in the limit of vanishing aspect ratio, as one can verify in Ref. [93]. The same plasma model, however, fails to predict the transverse oscillations of the chain since it becomes singular.
The short-wavelength spectrum is also amenable to some analytical treatment. In Ref. [224] the density of states at short wavelengths was determined by extending analytical methods, originally developed for calculating the dispersion relation of disordered chains [100]. The same treatment allows one to determine the largest eigenvalue $\lambda_{N}$. At leading order in an expansion in $1/\log N$ using that the rescaled term, $K_{0}\lambda_{N}$, is constant in the thermodynamic limit of Eq. (30), one obtains $K_{0}\lambda_{N}=(9/8)\nu^{2}N^{2}/\log N=8Q^{2}/(ma_{0}^{3})=8\omega_{0}^{2}$. This shows that the frequency $\omega_{0}$ scales the bandwidth of the axial mode spectrum. It also scales the critical value of the transverse trap frequency, $\nu_{t}^{(c)}$, at which the chain becomes unstable. The latter is found setting $\omega^{\perp}_{N}=0$, giving $\nu_{t}^{(c)}=\sqrt{K_{0}\lambda_{N}/2}=2\omega_{0}$. The corresponding critical aspect ratio reads $\alpha^{(c)}=(\nu/\nu_{t}^{(c)})=\sqrt{(16/9)\log N}/N$ [93, 278, 224]. In Sec. II.1.4 we review the structural transition at $\nu_{t}=\nu_{t}^{(c)}$ and its statistical properties.
@@
When instead the trap frequency $\nu_{t}\gg 8\omega_{0}$, there is a gap between axial and transverse modes. Moreover, the transverse spectrum becomes almost flat, so that the transverse dynamics is well approximated by local oscillators weakly coupled by the Coulomb interactions. This regime is at the center of the proposals for realizing effective Hubbard models [253] and for implementing quantum information processors [288] using ion chains.
@@ II.1.3 Thermodynamics of the ion chain
The thermodynamic properties of Debye crystals are textbook examples, and a linear chain provides a laboratory to verify hypotheses and predictions down to the quantum regime and close to the absolute zero. The behavior deep in the quantum regime can be accessed in ion chains thanks to laser cooling. Quantities such as the specific heat could be measured by a calorimetric experiment, where the temperature of the crystal is determined as a function of the energy transferred to the chain. The temperature can be extracted by resonance fluorescence, as in the protocols discussed in [268, 114], or by means of the so-called sideband thermometry [106, 327].
From the theoretical point of view, there is an essential difference between the textbook solid state crystals, where the interaction is typically nearest-neighbor, and the long-range Coulomb repulsion of ion Coulomb chains. In Ref. [225, 224] the specific heat per particle and the coefficient of thermal expansion have been determined for an ion chain in a steep transverse trap potential, assuming temperatures for which the transverse vibrations are frozen out. Using the notation of the previous section, this requires that $\nu_{t}\gg 8\omega_{0},k_{\mathrm{B}}T/\hbar$. In this regime the motion is effectively one dimensional.
The free energy is evaluated assuming quantized vibrations and a canonical ensemble with temperature $T$, using the scaling of Eq. (30), which warrants a well defined thermodynamic limit. The free energy is a function of the number of ions $N$, of the temperature $T$, and of the volume. The latter is determined by the length $L$ from end to end, which, for fixed $N$, is controlled by the axial trap frequency $\nu$, see Eq. (29). The specific heat is $c_{a}=(1/N)\partial U_{\rm th}/\partial T|_{\nu,N}$, where $U_{\rm th}$ is the thermal energy of a quantum gas of phonons, $U_{\rm th}=\sum_{n}\hbar\omega_{n}^{\|}/({\rm e}^{\beta\hbar\omega_{n}^{\|}}-1)$ with $\beta=(k_{B}T)^{-1}$ the inverse temperature. Figure 12(a) shows that the behavior of the specific heat with the temperature is very similar to the one of a one-dimensional Debye chain: it saturates at large temperatures, characteristic of Dulong-Petit law, while at low temperatures $c_{a}$ scales linearly with $T$. Interestingly, the limit $T\to 0$ and $N\to\infty$ do not commute. By keeping $N$ constant, the proportionality coefficient in the relation $c_{a}=\gamma T$ is a function of $N$, namely, $\gamma\sim N/\sqrt{\log N}$ [224]. This deviation from extensivity is due to the long-range interactions and becomes evident in the quantum regime. Despite the non-intensive behavior of the heat capacity, the usual ensemble equivalence holds, because the relative energy fluctuations at low temperatures scale as $(\sqrt{\log N}/N)^{1/2}$, and hence vanish in the thermodynamic limit $N\to\infty$.
@@
where $\kappa_{T}$ is the isothermal compressibility and is independent of $T$ at leading order [225]. Figure 12(b) shows that the functional dependence on the temperature is similar to the specific heat. Due to the non-intensive behavior of the isothermal compressibility, the coefficient of thermal expansion is a non-intensive function of $N$ over the whole range of temperatures.
The non-intensive behavior of specific heat and of $\alpha_{T}$ is a manifestation of the long-range, Coulomb interactions in one dimension. It is typically cured by theoretical procedures, so-called Kac’s scaling [45], which in this case would consist of artificially assuming that the ions electric charge depends on $N$ in such a way to re-establish extensivity of the energy [186]. In an experiment, where the charge is fixed, measurements at low temperatures would detect a relatively strong dependence of $c_{a}$ on the size of the chain, while they shall converge to a well defined, size-independent limit at relatively large temperatures.

### v2

@@ I.4 Ion trap arrays and optical dipole traps
In this review we focus on ion crystals that form at low temperatures in a configuration resulting from the interplay of a common ion trap potential and the Coulomb repulsion. In this section we provide a short outlook on other kinds of trapping potentials that have been discussed in the literature and are based on different mechanisms than the ones discussed in the rest of this review. One such alternative approach consists of using an array of trapping potentials, where each cell stores a single ion [319]. This approach enables the generation of arbitrary lattice geometries. Ion trap arrays are being pursued both with rf ion traps [236, 219, 260] as well as Penning traps [169, 170]. To date the number of ions stored in an ion trap array is small (up to $N\sim 4$).
A further approach employs optical dipole traps for ions with optical dipole transitions. A major motivation is to avoid boundary effects that are significant in Penning and Paul traps, thereby ideally realizing strings of equidistant ions. The micromotion due to the coupling of the charge monopole to the fast laser oscillations at the optical frequency is negligible and the potential the ions experience is essentially the same as the potential experienced by neutral atom scattering in the visible [69]. Optical trapping of ion Coulomb crystals is challenging because of the strength of the Coulomb repulsion of closely spaced ions, requiring large laser powers. Experimental demonstrations include trapping linear strings of up to 6 ions by a single beam optical dipole trap [316], and confinement of a few ions in a one-dimensional optical lattice [153]. An alternative approach is to use an array of optical tweezers [227]. Recent studies indicate that this architecture could provide a scalable approach for fault-tolerant digital quantum computation and for analog quantum simulations with ions arrays [7, 227, 322].
Somewhat less demanding in laser power is the use of optical dipole forces to modify the normal mode spectrum of trapped ion crystal. We refer the reader to Sec. III.5 for an extensive discussion.
@@ II Ion Coulomb crystals: from one to three dimensions
Trapped ions can crystallize in different structures, determined by the geometry of the confining potential and by their number. Linear, planar, up to cubic-centered structures have been realized in the laboratory. In this section we review the equilibrium properties of the crystals from one- to three-dimensional spatial arrangements of the ions. We start with the linear chain, on which an extensive number of studies have been performed, review then planar structures, and finally report on the crystalline properties of three-dimensional structures. Differing from previous reviews, we set the focus on the dynamical and thermodynamical properties, for the purpose of establishing a direct link with condensed-matter systems.
@@ II.1 Linear geometries: the ion chain
Figure 8 is the image of an experimentally realized ion chain composed of more than 100 ions [269]. The interparticle distances are of the order of several micrometers, and the individual ions can be imaged on a CCD camera via fluorescence light emitted during laser excitation. Ion chains typically form at Doppler cooling temperatures. First experimental demonstrations were in quadrupole traps [25, 291, 364]. In the meantime linear chains have also been realized in Penning traps [226] and ring geometries [209]. The ion chain has been the workhorse of the ion-trap quantum computer for the versatility in controlling the individual components, such as the interactions between the electronic and the vibrational degrees of freedom at the quantum level [64].
Chains of ions are a laboratory realization of a one-dimensional Debye crystal [8]. Due to the unscreened Coulomb interactions, it constitutes a peculiar one-dimensional structure. Below we review salient dynamic and thermodynamic properties, emphasizing analogies with and fundamental differences from textbook examples of condensed matter systems.
Figure 8: (Color online) Linear chain of 121 171Yb+ ions in a cryogenic linear Paul trap with $\omega_{y}/2\pi=1.5$ MHz and $\omega_{x}/2\pi=35$ kHz. The red arrow marks a lone 171Yb+ ion which is in a state that does not fluoresce. This dark ion retained its position after more than 4 hours, suggesting that the chain kept its ordered structure and never melted despite collisions with the background gas. From 269.
@@ II.1.1 The ground state configuration
where $m$ is the mass, $Q$ the charge, and $\nu$ and $\nu_{t}$ are respectively the trap axial frequency and the transverse trap frequencies in the pseudopotential approximation, such that $\nu<\nu_{t}$. The linear chain is stable for $\nu_{t}/\nu$ exceeding a critical value, as we will detail below.
@@
with $|x|\leq L$, and vanishes otherwise. The parameter $L$ is the chain half-length at equilibrium and is determined by means of a variational ansatz, such that to leading order in an expansion in $\log N$ it reads [101]
$L(N)=\left[N\log N(3Q^{2}/(m\nu^{2})\right]^{1/3}\,.$ (29)
At the chain center the density varies very slowly about the value $n(0)\sim N/L$, where the interparticle distance is minimal, $a(0)=1/n(0)$. Figure 9 shows that expression (28) gives a good estimate of the charge distribution sufficiently far away from the chain edges. This expression has been experimentally validated for chains composed of more than 100 ions and has been used to determine the number of trapped ions without counting the individual components [180].
Figure 9: Density of charges per unit length, $n(x)$, as a function of position $x$ from the center of the chain for different numbers $N$ of ions in a linear Paul trap. The dots are the equilibrium positions determined by molecular dynamics simulation for a finite number of ions $N$. The lines are obtained using the method of trial variational functions with two parameters. The dashed lines are the one-parameter function of Eq. (28). Both $x$ and $n(x)$ are scaled to $(Q^{2}/m\nu^{2})^{1/3}$. From 101.
@@
The above analysis ignores thermal and quantum fluctuations, which tend to suppress long-range order in one dimension. Strictly speaking, a perfectly ordered linear chain is unstable in the thermodynamic limit ($L\rightarrow\infty$). In practice, however, under suitable conditions the crystalline order can persist over sufficiently large length scales so that the chain effectively appears as fully ordered. Stability against quantum fluctuations has been studied in [321] by means of a treament based on bosonization, showing that at zero temperature the density-density correlations of a one-dimensional Wigner crystal decay with the distance $x$ as $\exp(-c\sqrt{\log x})$ with $c$ a constant [321]. This is an extremely slow decay: any large (but finite) chain can be considered crystallized. Stability against thermal fluctuations requires $Q^{2}\log N/(4\pi\epsilon_{0}a_{0})\gg\kappa_{B}T$, namely, the potential energy scale is larger than the thermal fluctuations where $T$ is the temperature and the classical equipartition theorem has been used [250]. This shows that, for sufficiently large chains, the long-range Coulomb correlations dominate over thermal disorder.
II.1.2 Normal modes of the ion chain
Figure 10: (Color online) The frequency sideband spectrum of the axial modes of a laser cooled chain composed of 24 9Be+ ions. Dashed lines are the experimental results of normalized excitation level versus frequency, where red and blue denote the scan over red-shifted sideband (anti-Stokes), carrier, and blue-shifted sideband (Stokes) transitions, respectively. The brightness of the dashed line shows the variable probe times throughout the spectrum to ensure that all modes are excited with high visibilities. Dashed vertical blue and red (dark) lines show all 24 calculated first-order red and blue sideband frequencies. Dashed vertical green (light) lines show the mode mixings between center-of-mass and a few higher frequency modes, coinciding with the smaller peaks in the experimental data. Details on the figures can be found in Ref. 372.
At laser cooling temperatures, the ions perform vibrations about the equilibrium positions with displacements of the order of tens of nanometers, and thus about three orders of magnitude smaller than the equilibrium interparticle distance. In this regime the restoring forces are harmonic and the ion chain is a one-dimensional harmonic crystal. Figure 10 displays the Stokes sidebands, measured in the resonance fluorescence of a chain of 24 ions. The measured spectra are reproduced by theoretical models that approximate the potential of Eq. (27) by only keeping the quadratic terms in the displacements about the equilibrium position.
Below we focus on ion chains in a linear Paul trap and in the pseudopotential approximation. We note that the analysis will also apply to linear chains in Penning traps [226] at the so-called Brillouin limit, where the effective magnetic field is zero. We refer to Refs. [214, 156, 52] for studies of the normal modes of ion chains in anharmonic traps.
@@ II.1.2 Normal modes of the ion chain
with $\delta_{ij}$ Kronecker’s delta. Here, $\xi_{i}=x_{i}^{(0)}/L$ is the dimensionless equilibrium position rescaled by the half-length $L$ of the chain at equilibrium. The set of differential equations (32) and (33) is solved by diagonalizing the matrix $I_{ij}$, thus solving the linear set of equations
@@
$\omega^{\perp\,2}_{n}=\nu_{t}^{2}-K_{0}\lambda_{n}/2\,$,
@@
In the rest of this section we make general considerations on the structure of the eigenvalues for $\nu_{t}>\nu_{t}^{(c)}$, which are based on the treatment of Ref. [249, 250]. For convenience, we adopt the convention $\lambda_{n}<\lambda_{m}$ for $n<m$. With this convention the label $n$ directly gives the number of nodes $\ell=n-1$ of the corresponding eigenmode $w^{(n)}$. Since the position of the nodes is generally constrained to the position of the ions within the chain, the value of $\lambda_{n}$ generally depends on the number of ions $N$. There are two exceptions: (i) The eigenvalue $\lambda_{1}=0$, corresponding to the bulk eigenmodes where the chain oscillates as a rigid body at the frequencies of the trap, and (ii) the eigenvalue $\lambda_{2}=2\nu^{2}/K_{0}$, corresponding to the breathing mode, which has a single node at the chain center. These two eigenvalues are related to two symmetries of the Hamiltonian: point (i) is due to the fact that the dynamics is separable into center-of-mass and relative coordinates even in the presence of the harmonic trap, point (ii) is a consequence of the symmetry by reflection about the trap center. The breathing oscillations $\omega_{2}^{\|}=\sqrt{3}\nu$ and $\omega_{2}^{\perp}=\sqrt{\nu_{t}^{2}-\nu^{2}}$, in fact, are also characteristic of a one-component plasma in an ellipsoidal trap [106].
Figure 11: (a) Transverse and (b) axial spectrum of eigenfrequencies (in units of $\nu$) for a chain of $N=1000$ ions and with $\nu_{t}=400\nu$. The solid line is obtained from the numerical solution of Eq. (35), the grey line from the spectrum of Eq. (38) (the grey curves have been truncated, as they do not correctly reproduce the shortwavelength eigenmodes). The dotted line gives the spectra evaluated extending a method first developed for determining the density of states of a disordered chain. Adapted from Ref. 250.
@@
The eigenmodes are Jacobi polynomials $w^{(n)}=P_{\ell}^{(1,1)}$ with $n=\ell+1$ and $\ell=0,1,\ldots$ The eigenvalues read ${\lambda}_{n}=\ell(\ell+3)$ and exactly reproduce the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ for $\ell=0,1$, as also visible in Fig. 11. Interestingly, the model of a plasma in a cigar-shaped trap potential gives the same expression for the axial spectrum $\omega_{n}^{\|}=\nu\sqrt{n(n+1)/2}$ in the limit of vanishing aspect ratio, see Ref. [99]. The same plasma model, however, fails to predict the transverse oscillations of the chain since it becomes singular.
The short-wavelength spectrum is also amenable to some analytical treatment. In Ref. [250] the density of states at short wavelengths was determined by extending analytical methods, originally developed for calculating the dispersion relation of disordered chains [107]. The same treatment allows one to determine the largest eigenvalue $\lambda_{N}$. At leading order in an expansion in $1/\log N$ and using the thermodynamic limit of Eq. (30), one obtains $K_{0}\lambda_{N}=(9/8)\nu^{2}N^{2}/\log N=8Q^{2}/(ma_{0}^{3})=8\omega_{0}^{2}$. This shows that the frequency $\omega_{0}$ scales the bandwidth of the axial mode spectrum. It also scales the critical value of the transverse trap frequency, $\nu_{t}^{(c)}$, at which the chain becomes unstable. The latter is found setting $\omega^{\perp}_{N}=0$, giving $\nu_{t}^{(c)}=\sqrt{K_{0}\lambda_{N}/2}=2\omega_{0}$. The corresponding critical aspect ratio reads $\alpha^{(c)}=(\nu/\nu_{t}^{(c)})=\sqrt{(16/9)\log N}/N$ [99, 311, 250]. In Sec. II.1.4 we review the structural transition at $\nu_{t}=\nu_{t}^{(c)}$ and its statistical properties.
@@
When instead the trap frequency $\nu_{t}\gg 8\omega_{0}$, there is a gap between axial and transverse modes. Moreover, the transverse spectrum becomes almost flat, so that the transverse dynamics is well approximated by local oscillators weakly coupled by the Coulomb interactions. This regime is at the basis of the proposals for realizing effective Hubbard models [282] and for implementing quantum information processors [323] using ion chains.
@@ II.1.3 Thermodynamics of the ion chain
The thermodynamic properties of Debye crystals are textbook examples, and a linear chain provides a laboratory to verify hypotheses and predictions down to the quantum regime and even close to the absolute zero, thanks to laser cooling. Quantities such as the specific heat could be measured by a calorimetric experiment, where the temperature of the crystal is determined as a function of the energy transferred to the chain. The temperature can be extracted by resonance fluorescence, as in the protocols discussed in [300, 123], or by means of the so-called sideband thermometry [113, 363].
From the theoretical point of view, there is an essential difference between the textbook solid state crystals, where the interaction is typically nearest-neighbor, and the long-range Coulomb repulsion of ion Coulomb chains. In Ref. [249, 250] the influence of the Coulomb repulsion on the thermodynamic behavior has been studied by determining the specific heat per particle and the coefficient of thermal expansion. This work considered the case in which the motion is effectively one-dimensional, which is realized for an ion chain in a steep transverse trap potential, assuming temperatures for which the transverse vibrations are frozen out. Using the notation of the previous section, this requires that $\nu_{t}\gg 8\omega_{0},k_{\mathrm{B}}T/\hbar$.
The free energy is evaluated assuming quantized vibrations and a canonical ensemble with temperature $T$, using the scaling of Eq. (30), which warrants a well defined thermodynamic limit. The free energy is a function of the number of ions $N$, of the temperature $T$, and of the volume. The latter is determined by the length $L$ from end to end, which, for fixed $N$, is controlled by the axial trap frequency $\nu$, see Eq. (29). The specific heat is $c_{a}=(1/N)\partial U_{\rm th}/\partial T|_{\nu,N}$, where $U_{\rm th}$ is the thermal energy of a quantum gas of phonons, $U_{\rm th}=\sum_{n}\hbar\omega_{n}^{\|}/({\rm e}^{\beta\hbar\omega_{n}^{\|}}-1)$ with $\beta=(k_{B}T)^{-1}$ the inverse temperature. Figure 12(a) shows that the behavior of the specific heat with the temperature is very similar to the one of a one-dimensional Debye chain: it saturates at large temperatures, characteristic of Dulong-Petit law, while at low temperatures $c_{a}$ scales linearly with $T$. Interestingly, the limit $T\to 0$ and $N\to\infty$ do not commute. By keeping $N$ constant, the proportionality coefficient in the relation $c_{a}=\gamma T$ is a function of $N$, namely, $\gamma\sim N/\sqrt{\log N}$ [250]. This deviation from extensivity is due to the long-range interactions and becomes evident in the quantum regime. Despite the non-intensive behavior of the specific heat (namely, the fact that it depends on $N$), the usual ensemble equivalence holds, because the relative energy fluctuations at low temperatures scale as $(\sqrt{\log N}/N)^{1/2}$, and hence vanish in the thermodynamic limit $N\to\infty$.
@@
where $\kappa_{T}$ is the isothermal compressibility and is independent of $T$ at leading order [249]. Figure 12(b) shows that the functional dependence on the temperature is similar to the specific heat. Due to the non-intensive behavior of the isothermal compressibility, the coefficient of thermal expansion is also a non-intensive function of $N$ over the whole range of temperatures.
The non-intensive behavior of specific heat and of $\alpha_{T}$ is a manifestation of the long-range, Coulomb interactions in one dimension. It is typically cured by theoretical procedures, so-called Kac’s scaling [49], which in this case would consist of artificially assuming that the ions electric charge depends on $N$ in such a way to re-establish extensivity of the energy [197]. In experiments the charge is fixed and measurements at low temperatures would detect a relatively strong dependence of $c_{a}$ on the size of the chain, while they shall converge to a well defined, size-independent limit at relatively large temperatures.